RobertRobert

is buying

DVDsDVDs.

He receives

8888

utils from his first

DVDDVD​,

4444

utils from his second

DVDDVD​,

and

2727

utils from his third

DVDDVD.

​(Round your answers to the nearest​ integer.)

RobertRobert​'s

marginal utility from his third

DVDDVD

is

nothing.

RobertRobert​'s

total utility from consuming three

DVDsDVDs

is

nothing.
Answer #1

Total utility received from purchase of first DVD = 88

Total utility received from the purchase of second DVD = 44 + 88 = 132

Total utility received from the purchase of third DVD = 27 + 132 = 159

Marginal utility is the additional utility received from the purchase of one more unit. Marginal utility from third DVD = total utility from third DVD - total utility from second DVD

= 159 - 132 = 27

Therefore marginal utility from third DVD is 27

Total utility from consuming 3 DVDs is 159